OpenStudy (anonymous):

[Lab 8] Can anyone tell me how I found the 'Component value for Device #1' in Circuit 2 and 3?

OpenStudy (anonymous):

The answers are very dependent on the curves you have

OpenStudy (anonymous):

The easiest way to determine what the 2 components is is to draw it in the sandbox or you can just use intuition since we know what R, L and C plot look like. (remember that at least one of the components must be a 1K ohm resistor,) to find the correct value you have to determine if it is rising or falling then use the right equation: C: Vf= Vi(1-e^-t/RC) or Vf= Vi(e^-t/RC) L: Vf= Vi(1-e^-t*R/L) or Vf= Vi(e^-t*R/L) Voltage divider:V*(R2/R1+R2) And just solve with the value from the graph. *Hint: for rise I used Vi=1 hope that helps. Good luck